OpenStudy (abmon98):

Integrate e^3x(1+e^2x)^1/2

zepdrix (zepdrix):

Here is one approach, \[\large\rm \int\limits e^{3x}\sqrt{1+e^{2x}}~dx\quad=\int\limits \left(e^x\right)^2\sqrt{1+\left(e^x\right)^2}~e^x~dx\]From there you can make the substitution,\[\large\rm \tan \theta= e^x\]

zepdrix (zepdrix):

Works out well enough, you have to apply Secant Reduction Formula at one point.

zepdrix (zepdrix):

There's probably some way to do Integration by Parts, I'm just not seeing it.

OpenStudy (abmon98):

I tried integration by parts but you would keep on getting increasing power of e^x.

OpenStudy (anonymous):

Have you considered using substitution, two times. First substitute \[u=e^x\]Then substitute \[u=tan(s)\]

OpenStudy (anonymous):

And then you get something where you need to use the reduction formula: \[\int sec^m(s)ds=\frac{sin(s)sec^{m-1}(s)}{m-1}+\frac{m-2}{m-1}\int sec^{-2+m}(s)ds\]

OpenStudy (abmon98):

@zepdrix I never studied reduction formulas before i don't get why we choose our substitution for e^x to be tanx. I used your substitution and it worked well

OpenStudy (solomonzelman):

Are you sure that the first thing is \(e^{3x}\), and not \(e^{2x}\)? That would be indeed much easier:)

OpenStudy (irishboy123):

here you have \(\sqrt {1 + blah^2}\), and so you want to sort that bit out right away the identity \(\tan^2 x + 1 = \sec^2 x\) is prolly the most obvious way to get rid of that radical

OpenStudy (irishboy123):

so if it was \(\sqrt{1- e^{2x}}\) you'd look at the cos sin identity

OpenStudy (abmon98):

\[\int\limits_{}^{}\tan x \tan x^2\sqrt{secx^2}dx=\int\limits_{}^{}\tan x(\sec^2x-1)secxdx=\int\limits_{}^{}\tan xsec^3xdx-\int\limits_{}^{}tanxsecx\]

OpenStudy (abmon98):

i got my answer as secx^3/3-secx

OpenStudy (irishboy123):

it should come out as \(\int \; \sec^5 \theta - \sec^3 \theta \; d\theta\) then those reductions things kick in

zepdrix (zepdrix):

Ya I think you made a boo boo somewhere, hmm..\[\large\rm \color{royalblue}{e^x=\tan \theta},\qquad\qquad\qquad \color{orangered}{e^x~dx=\sec^2\theta~d \theta}\]Gives us,\[\rm \int\limits\limits (\color{royalblue}{e^x})^2\sqrt{1+(\color{royalblue}{e^x})^2}~\color{orangered}{~e^x~dx}\qquad=\qquad\int\limits\limits (\color{royalblue}{\tan \theta})^2\sqrt{1+(\color{royalblue}{\tan \theta})^2}~\color{orangered}{\sec^2 \theta~d \theta}\]Which simplifies to,\[\large\rm \int\limits \tan^2 \theta \sec^3 \theta~d \theta\]I don't think you ended up with enough secants in yours.

OpenStudy (solomonzelman):

I know it is not the case, by why isn't like this? \(\color{#000000 }{ \displaystyle \sqrt{\sec^2\theta}=\color{red}{\left|\color{black}{\sec \theta}\right|} }\) ? That would (but doesn't) go according to the definition of the absolute value, i.e. \(\color{#000000 }{ \displaystyle \sqrt{\varphi^2}=\left|\varphi \right| }\)

OpenStudy (irishboy123):

@SolomonZelman i agree, i see that all the time with trig subs

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \sqrt{1+\tan^2\theta~}=\sqrt{\sec^2\theta}=\color{red}{\left|\color{black}{\sec \theta}\right|} }\) maybe something ehre that makes sec\(\theta\) undefined for negatives(?)

zepdrix (zepdrix):

Well I remember when doing First-Order ODE's, for integrating factor you didn't have to be too picky,\[\large\rm \mu=e^{\int\limits p(x)dx}\]We didn't have to be careful and do something like this,\[\large\rm e^{\int\limits\limits 2x~dx}\quad=e^{x^2+c}\]because we were looking for `a suitable substitution`, not a family of substitutions or anything. We just needed one that worked. We just didn't need the +c. Maybe that is what's going on here? We just care about the principal root because ummmm Hmm no -_- hmm

zepdrix (zepdrix):

Err not substitution :) But integrating factor* blah, whatever, i failed to make any kind of point there anyway lol

OpenStudy (anonymous):

Refer to the tiff attachment.

zepdrix (zepdrix):

Woops rob :( The x's be in the exponent, e^{3x} not e^3x

OpenStudy (anonymous):

A detailed solution from WolframAlpha through Mathematica v9 is attached.